Let ??~N(3, 9) and ?? = 5 ? ??.

a) Find P(X > 2)

b) Find P(?1 < Y < 3)

c) Find P(X > 4|Y < 2)

Please use an approximation to calculate the value of ?(x), where necessary
Answer #1

Here X~N(3,9) and Y=5-X

a)

P(X>2)=P(Z>\frac{2-3}{\sqrt{9}})=P(Z>-0.33)=P(Z<0.33)=0.6293

b)

P(-1<Y<3)=P(-1<5-X<3)=P(-6<-X<-2)

=P(6>X>2)=P(2<X<6)=P(\frac{2-3}{\sqrt{9}}<Z<\frac{6-3}{\sqrt{9}})

=P(-0.33<Z<1)=P(Z<1)-P(Z<-0.33)

=P(Z<1)-P(Z>0.33)=P(Z<1)-[1-P(Z<0.33)]

=0.8413-[1-0.6293]=0.4706

c)

P(X>4|Y<2)=\frac{P(X>4\cap Y<2)}{P(Y<2)}=\frac{P(X>4\cap 5-X<2)}{P(5-X<2)}

=\frac{P(X>4\cap X>3)}{P(X>3)}=\frac{P(X>4)}{P(X>3)}=\frac{P(Z>\frac{4-3}{\sqrt{9}})}{P(Z>\frac{3-3}{\sqrt{9}})}=\frac{P(Z>0.33)}{P(Z>0)}

=\frac{1-P(Z<0.33)}{1-P(Z<0)}=\frac{1-0.6293}{1-0.5}=0.7414